Please help me with this question: You have 32 coins that equals $6.20. How many dimes and quarters do you have ?
Answer by jim_thompson5910(35256)

Hint:

"You have 32 coins" ----> d+q = 32 which becomes d = 32 - q


"...that equals $6.20" ----> 0.1d + 0.25q = 6.20 which becomes 10d + 25q = 620

-------------------------------------------------------

Now plug d = 32 - q into the last equation

10d + 25q = 620

10(32 - q) + 25q = 620

and solve for q. Then use this solution to find d.